OpenStudy (anonymous):

The functions s and t are defined as follows. s(x)= 2x-1 t(x)= -x^2-1 Find the value of s(t(4)).

OpenStudy (jwang55):

=2(-(4)^2-1)-1 =2(16-1)-1 =2(15)-1 =30-1 =29